### v2

Abstract
Given the low throughput of blockchains like Bitcoin and Ethereum, scalability — the ability to process an increasing number of transactions — has become a central focus of blockchain research. One promising approach is the parallelization of transaction execution across multiple threads. However, achieving efficient parallelization requires a redesign of the incentive structure within the fee market. Currently, the fee market does not differentiate between transactions that access multiple high-demand storage keys (i.e., unique identifiers for individual data entries) versus a single low-demand one, as long as they require the same computational effort. Addressing this discrepancy is crucial for enabling more effective parallel execution.
In this work, we aim to bridge the gap between the current fee market and the need for parallel execution by exploring alternative fee market designs. To this end, we propose a framework consisting of two key components: a Gas Computation Mechanism (GCM), which quantifies the load a transaction places on the network in terms of parallelization and computation, measured in units of gas, and a Transaction Fee Mechanism (TFM), which assigns a price to each unit of gas. We additionally introduce a set of desirable properties for a GCM, propose several candidate mechanisms, and evaluate them against these criteria. Our analysis highlights two strong candidates: the weighted area GCM, which integrates smoothly with existing TFMs such as EIP-1559 and satisfies a broad subset of the outlined properties, and the time-proportional makespan GCM, which assigns gas costs based on the context of the entire block’s schedule and, through this dependence on the overall execution outcome, captures the dynamics of parallel execution more accurately.
keywords:
blockchain, transaction fee mechanism, parallel execution
category:
1 Introduction
Scalability, the ability to process more transactions efficiently, has become a central focus in blockchain research, especially given the low throughput of many existing networks. Ethereum, for example, is constrained by its single-threaded execution model, limiting transaction throughput. One promising way to enhance scalability is by parallelizing transaction execution across multiple threads, taking advantage of the multi-core processors common in modern hardware. However, achieving the full efficiency gains of parallel execution requires rethinking the fee market design to better account for storage key contention and scheduling constraints. A storage key is a unique identifier (like an address label) for a specific data item in storage.
A transaction fee mechanism (TFM) is a core component of any blockchain protocol, determining which pending transactions should be processed and what users must pay for the privilege of having their transactions executed. Traditional fee mechanisms, like Bitcoin’s first-price auction, involve users submitting bids with their transactions, and the transactions with the highest bids per computation are included in the next block. Ethereum initially used a similar purely first-price auction-based model but switched to the more sophisticated EIP-1559 mechanism in 2021, which introduces a fluctuating base fee based on network demand, aiming to improve incentive compatibility and reduce price volatility [11]. Most of the existing literature on transaction fee mechanisms focuses on settings where transactions are executed sequentially and therefore does not account for storage key contention, which is crucial in parallel execution environments.
Thus, viewed in isolation, these traditional TFMs are ill-suited to the complexities introduced by parallel transaction execution. They price transactions solely based on their computational cost, without distinguishing between those that access a single storage key and those that interact with multiple, potentially contested storage keys. This pricing model works in a single-threaded environment but fails to capture the nuances of parallel execution, where transactions may impose vastly different constraints on storage key scheduling. Transactions that touch multiple high-contention storage keys can introduce bottlenecks, while those interacting with isolated storage keys are far easier to schedule efficiently.
Ethereum has yet to adopt parallel execution, but several blockchains, such as Solana, Aptos, and Sui, already employ parallel transaction processing [4, 42, 56, 59, 45]. However, many of these networks have yet to implement fee models that fully account for the challenges of parallel execution. Sui and Solana have introduced fee markets tailored to parallelization, but these mechanisms require users to engage in first-price auctions for congested storage keys [39, 43]. As a result, these fee markets demand a high level of sophistication from users to effectively optimize their fee settings and are also not incentive-compatible. The requirements for an effective fee market that is suitable for parallel execution and the design of such a market have so far remained unresolved.
Our Contributions.
In this work, we aim to bridge this gap by outlining the requirements for such a fee market and evaluating possible candidates. We outline our main contributions below:
  • We introduce a framework with two main components: a Gas Computation Mechanism (GCM), which measures the load a transaction imposes on the network in terms of both parallel execution and computation, expressed in units of gas, and a Transaction Fee Mechanism (TFM), which determines the cost associated with each unit of gas.
  • We introduce a list of desirable properties for a GCM and evaluate against them a set of mechanisms that we propose.
  • Our analysis identifies two promising candidates. The weighted area GCM achieves a large subset of the outlined properties and, importantly, can be seamlessly integrated with existing TFMs, inheriting their properties. Complementing it, the time-proportional makespan GCM is designed to price execution costs in proportion to the total computational load of a block, allowing more accurate resource pricing at the block level.
2 Model
We consider a universe consisting of several stateful storage keys (e.g., user accounts, storage addresses of smart contracts). Each storage key can be thought of as a system global variable. We write $\mathcal{K}$ for the set of storage keys. For analysis purposes, we assume that $\mathcal{K}$ is infinite. A transaction is a sequence of elementary instructions performing computation and interacting with the storage keys. Some of these operations access a given target storage key (e.g., read its value, write to it). For simplicity, we assume that the following are known in advance and supplied with the transaction:
  • The non-empty set of storage keys $K\subseteq\mathcal{K}$ the transaction accesses, or an overestimate. [^2] In Ethereum, transaction accesses are generally not known in advance. Instead, transactions can execute arbitrary logic (constrained by a maximum amount of computational effort) and their execution depends on the blockchain’s state at the time of execution. Ethereum currently supports optional access lists that allow transactions to specify their accesses [13]. This optional list could be made mandatory to provide an overestimate of accesses. Additionally, it is worth noting that in other blockchains (e.g., Solana [35] and Sui [34]), an overestimate of accesses is typically known in advance.
  • The total time $t>0$ it takes to execute the transaction on a single thread. [^3] In reality, the time to execute a transaction depends on the hardware it is executed on. Thus, time is estimated in Ethereum by assigning each operation a computational effort. Then the sum of the computational effort of a transaction’s operation can be seen as a proxy for time.
Transactions execute concurrently, but atomically, meaning that the overall effect of executing a batch of transactions should also be achievable by a sequential, single-core execution. For simplicity, we restrict to concurrent schedules following a simple lock-based execution policy: each storage key has a lock associated with it; whenever a thread wants to execute a transaction, it first locks all required storage keys, then executes the transaction, and then releases the locks. We assume that acquiring (and similarly releasing) the required locks happens simultaneously and takes no additional time. These simplifications have a desirable side-effect: $t$ and $K$ for each transaction now uniquely determine the set of admissible concurrent schedules, allowing us to ignore other details about the transactions:
Definition 2.1 (Transaction).
A transaction tx is specified through a tuple $(t,K)$, where $t>0$ denotes the time required to execute the transaction and $K\neq\varnothing$ represents the set of storage keys the transaction demands, which are locked for the duration of the transaction. We will sometimes write $t(\textit{tx})$ and $K(\textit{tx})$ for $t$ and $K$ respectively.
To give the tuple associated with a transaction, we will write $\textit{tx}\simeq(t,R)$. Note that different transactions might have the same associated tuple. We will also write $\textit{tx}_{1}\simeq\textit{tx}_{2}$ to mean that the two transactions have the same associated tuple.
In Figure 1, we illustrate a sample transaction $\textit{tx}\simeq(3,\{k_{2},k_{3}\})$. Transaction tx thus takes 3 units of time to execute and utilizes storage keys $k_{2}$ and $k_{3}$. We illustrate this by a rectangle of corresponding length (i.e., time) and width (i.e., storage keys). Throughout, we will illustrate transactions in this manner to aid in visualizing concepts and results. Note that we will always represent transactions as rectangles (i.e., they use consecutive storage keys). In reality, this is, of course, not the case. Importantly, all our results also hold in the more general setting where a transaction can use any subset of storage keys.
[…]
Figure 1: Illustration of a sample transaction $\textit{tx}\simeq(3,\{k_{2},k_{3}\})$.
A blockchain is a sequence of blocks comprising bundles of transactions: $B_{1},\dots,B_{m}$. The system starts in some predetermined initial state. The blocks are executed in order, starting from the oldest (the genesis block $B_{1}$), successively changing the system’s state. For simplicity, we assume no cross-block parallelism, so the execution of a block only starts after the previous block’s execution has been completed. However, the execution of transactions inside a block happens concurrently. The system’s state after executing $B_{m}$ is the current system state. Users wanting to change the system’s state compete for inclusion in the next block $B_{m+1}$ and have to pay a fee if successfully included. Desirably, the fee should be higher for more complex transactions and higher during periods of high demand due to limited block space. These requirements are typically decoupled and ensured through different means:
  1. 1.
    A transaction’s complexity is quantified in units of gas: [^4] For Ethereum, the unit is called wei. the more complex a transaction is, the more gas it consumes. Gas encompasses multiple components such as execution, storage space, and data bandwidth. For our purposes, we will only be concerned with execution gas. [^5] From this point forward, gas will refer specifically to execution gas. Currently, the execution gas acts as a proxy for the execution time tt, [^6] Throughout we will assume this approximation to the exact. but this need not be the case, as we will demonstrate in our paper.
  2. 2.
    The fair competition for block space is ensured through a transaction fee mechanism (TFM): users submit transactions they would like to be included in the next block together with bids of what they would be willing to pay per unit of gas. Importantly, block space is limited, i.e., there is limited space for transactions. The mechanism then determines the set [^7] In practice, blocks are ordered, and transactions often compete for earlier positions (particularly in MEV settings). Following a common simplification in the TFM literature, we model blocks as sets of transactions. Accounting for intra-block ordering would require extending our model and constraining it to schedulers that honor the specified ordering. Our framework appears reasonably extensible to such settings, and we leave a detailed treatment to future work. of transactions to be included in the block, together with a price per unit of gas to be paid by each included transaction (potentially not the same for all transactions). [^8] There are several other components of a TFM, but for the level of detail we need here, this suffices.
As such, an included transaction consuming $g$ gas units at a price of $p$ per unit of gas will have to pay a fee of $g\cdot p$ (generally in the blockchain’s native currency).
To keep the separation of concerns in (1) and (2), we will keep the formula for the fee $g\cdot p$ and instead vary the gas computation mechanism used to determine $g$:
Definition 2.2 (Gas Computation Mechanism).
A gas computation mechanism (GCM) takes as input a set of transactions $T$ to be included in a block and determines in a deterministic [^9] We focus solely on deterministic mechanisms. This is common in blockchain fee markets, given the inherent difficulty of accessing true randomness on-chain (though notable approaches exist, such as VRFs). While randomized mechanisms can statistically improve over their deterministic counterparts, their ex-post behavior is often hard to predict, leading to higher fee uncertainty, an outcome generally undesirable in blockchain settings. manner the amount of gas consumed by each transaction $\textit{tx}\in T$, written $\textsf{gas}_{T}(\textit{tx})$.
Currently deployed GCMs associate a fixed, predetermined gas consumption with each transaction, independent of the specific storage keys accessed by the transaction, making them unsuitable for a parallel execution environment. In particular, this is true for Ethereum’s current GCM:
Definition 2.3 (Current GCM).
Given a set of transactions $T$ and a transaction $\textit{tx}\in T$ with $\textit{tx}\simeq(t,R)$, the current GCM computes the amount of gas used by tx as follows:
$\textsf{gas}_{T}^{C}(\textit{tx}):=t.$
Since this does not depend on $T$, we often drop the subscript.
Our goal is to ensure that fees accurately reflect the parallelizability of transactions. Therefore, the gas consumption calculated for a transaction should depend on the set of storage keys it accesses and may also be influenced by factors external to the transaction itself (like interactions between transactions).
3 GCM Properties
Next, we outline several desirable properties that a GCM should possess to provide the right incentives for parallelization. These properties should be viewed as a wishlist — as will become clear later on, no single mechanism can satisfy all of them simultaneously.
We begin with two natural monotonicity properties, one for storage keys and one for time. First, a transaction that requires a subset of the storage keys used by another transaction, while taking the same amount of time, should consume no more gas (Properties˜1 and 5). Similarly, a transaction that requires no more execution time than another, assuming both involve the same set of storage keys, should consume no more gas (Properties˜2 and 5).
[…]
Figure 2: Illustration of Property˜1, where $T=\emptyset$,    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$.
[…]
Figure 3: Illustration of Property˜2, where $T=\emptyset$,    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$.
[…]
Figure 4: Illustration of Property˜4, where $T=\emptyset$, $T_{1}=\{\textit{tx}_{1},tx_{2}\}$, and $T_{2}=T_{1}\cup\{\textit{tx}_{3}\}$. Here,    represents a sample $\textit{tx}_{1}$,    represents a sample $\textit{tx}_{2}$, and    represents a sample $\textit{tx}_{3}$.
[…]
Figure 5: Illustration of Property˜5, where $T=\emptyset$,    represents a sample $\textit{tx}_{3}$, while    represents a sample $\textit{tx}_{1}$ and    represents a sample $\textit{tx}_{2}$. The darker shaded areas indicate when a transaction operates on a storage key.
Property 1 (Storage Key Monotonicity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}=t_{2}$ and $K_{1}\subseteq K_{2}$:
$\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$
Property 2 (Time Monotonicity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}=K_{2}$:
$\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$
The previous two properties fix one dimension while varying the other. One can also define a seemingly stronger property that allows both to vary, as follows:
Property 3 (Storage Key-Time Monotonicity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$:
$\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$
Intuitively, if $\textit{tx}_{1}\lesssim\textit{tx}_{2}$, by which we mean $t(\textit{tx}_{1})\leq t(\textit{tx}_{2})$ and $K(\textit{tx}_{1})\subseteq K(\textit{tx}_{2})$, then $\textit{tx}_{1}$ should cost no less than $\textit{tx}_{2}$. However, an attentive reader will observe that the former two are collectively equivalent to the latter (the proof and all subsequent omitted proofs can be found in the appendix):
Lemma 3.1.
Property˜3 holds if and only if Properties˜1 and 2 hold.
Let us now take a moment to briefly evaluate why Properties˜1, 2 and 3 are not merely intuitive, but their violation can lead to harmful consequences and misaligned incentives: suppose $\textit{tx}_{1}\lesssim\textit{tx}_{2}$ and $T$ is a set of transactions containing neither of the two. If $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})>\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2})$, a user intending to submit $\textit{tx}_{1}$ might instead pad $\textit{tx}_{1}$ with unnecessary instructions and declare a larger access list to decrease the gas consumption. This would paradoxically reduce the gas usage and, assuming a reasonable TFM is used to compute transaction fees, also lower the transaction fee.
For any of the three properties above, we say that the property is strictly satisfied if for $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$ the conclusion inequality holds strictly. Note that properties holding strictly are even more desirable with respect to the reasoning above: replacing a transaction with a “larger” one is then not only no better but actively worse. Unsurprisingly, Lemma˜3.1 also holds for the strict versions of the properties:
Lemma 3.2.
Property˜3 holds strictly if and only if Properties˜1 and 2 hold strictly.
Next, we introduce another desirable monotonicity property, this time with a different emphasis: if two sets of transactions satisfy $T_{1}\subseteq T_{2}$, the transactions in $T_{1}$ should collectively consume no more gas than those in $T_{2}$ (Properties˜4 and 5). To formalize this, given a set of transactions $T$ and a subset $T^{\prime}\subseteq T$, write $\textsf{gas}_{T}(T^{\prime}):=\sum_{\textit{tx}^{\prime}\in T^{\prime}}\textsf{gas}_{T}(\textit{tx}^{\prime})$ for the total gas consumed by the transactions in $T^{\prime}$ when included in the set $T$ that constitutes a block.
Property 4 (Set Inclusion).
Given a set of transactions $T$ and two sets of transactions $T_{1}\subseteq T_{2}$, disjoint from $T$:
$\textsf{gas}_{T\cup T_{1}}(T_{1})\leq\textsf{gas}_{T\cup T_{2}}(T_{2}).$
To understand why this property is desirable, consider a GCM for which it does not hold. Then, there must be a scenario where a set of transactions can reduce their total gas consumption by adding additional transactions. This situation could be exploited through collusion by the users originating these transactions. Importantly, such a possibility is undesirable, as it would primarily benefit sophisticated users capable of orchestrating such arrangements.
Similarly to before, we say that Property˜4 holds strictly if for $T_{1}\neq T_{2}$ the inequality in the conclusion holds strictly, which is desirable for reasons similar to those discussed above.
We now move on to more complex properties. Given two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2}),$ a transaction $\textit{tx}_{3}$ is their sequential composition (or, more simply, concatenation), if it executes the steps of $\textit{tx}_{1}$ followed by the steps of $\textit{tx}_{2}$. Note that, in this case $\textit{tx}_{3}\simeq(t_{1}+t_{2},K_{1}\cup K_{2})$. Our next property states that the concatenation of two transactions should consume no less gas than submitting them individually (Properties˜5 and 5). The two individual transactions perform the same actions as their concatenation, but not atomically, with no guarantee of their relative ordering or control over what happens between them. Naturally, enforcing atomicity and ordering limits the set of admissible concurrent schedules and requires storage keys to remain locked over longer continuous timespans. In particular, $\textit{tx}_{3}$ requires the storage keys in $K_{1}\cup K_{2}$ to be locked over a continuous span of $t_{1}+t_{2}$ units, while $\textit{tx}_{1}$ and $\textit{tx}_{2}$ submitted individually only require exclusive access to storage keys in $K_{1}$ for $t_{1}$ units and to storage keys in $K_{2}$ for $t_{2}$ units. Hence, the “bigger” transaction is at least as hard to schedule as its two constituent “parts” and should hence consume no less gas.
Property 5 (Transaction Bundling).
Consider a set of transactions $T$ and three transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\notin T$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$, then:
$\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}(\textit{tx}_{1})+\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}(\textit{tx}_{2})\leq\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).$
Let us again consider the risks of having a GCM that does not satisfy the previous property. In such a scenario, a group of users could collude to collectively consume less gas by combining their transactions into a single transaction rather than processing them individually. This outcome would be undesirable, particularly because it disproportionately benefits sophisticated users, as we outlined before. Note that this could even be the case for a single user wanting to submit multiple transactions.
We say that Property˜5 holds strictly if the inequality in the conclusion holds strictly, which is again more desirable than the basic version of the property.
Next, we would like to formalize the intuitive idea that a transaction’s gas consumption fairly reflects its impact on the execution time. To do so, we first need to formalize scheduling more precisely. Let $n\geq 2$ be the number of available threads.
Definition 3.3.
A scheduler (for $n$ threads) takes as input a set of transactions $T$, and outputs a concurrent schedule using at most $n$ threads to execute all transactions in $T$. This schedule specifies the operations each thread should perform and the order in which they should be performed.
The following conditions should hold for any generated schedule:
  • Each transaction is assigned to a single thread;
  • A thread can work on only one transaction at a time;
  • There is no preemption: once a thread starts executing a transaction, it completes the transaction without context switching;
  • Transactions with overlapping storage key access sets cannot be executed in parallel: one must finish before the other can begin.
[…]
Figure 6: Illustration of optimal schedules for a set of four transactions: $\textit{tx}_{1}\simeq(2,\{k_{2},\dots,k_{8}\})$ (shown in   ), $\textit{tx}_{2}\simeq(4,\{k_{2},k_{3}\})$ (shown in   ), $\textit{tx}_{3}\simeq(5,\{k_{4},k_{5},k_{6}\})$ (shown in   ), and $\textit{tx}_{4}\simeq(2,\{k_{7},k_{8}\})$ (shown in   ). In the left plot, we show the optimal schedule for $n=3$, and in the right plot for $n=2$. Notice how for $n=2$, we cannot schedule the three transactions $\textit{tx}_{2}$, $\textit{tx}_{3}$, and $\textit{tx}_{4}$ to be executed in parallel even though they access pairwise-disjoint sets of storage keys.
Note that some of these conditions are natural for scheduling in general, while others arise from us assuming the simple lock-based execution policy. [^10] Using locks is one way to enforce this policy, but in our case — where the contents of a block are known before execution commences — it can also be achieved without locks, as long as the execution environment ensures that threads strictly follow a pre-determined schedule. We chose this name because it is largely suggestive of the intended semantics. For our purposes, we are not concerned with which thread executes which transaction, but only that no more than $n$ transactions ever execute simultaneously (so we can draw concurrent schedules as in Figure˜6, which illustrates the concept). To determine gas consumption, our notation will need to capture even less — given a scheduler $s$ (for $n$ threads), we write $v^{s}(T)$ for the makespan of the schedule produced by $s$ for the set of transactions $T$ (i.e., the time required to execute the schedule in parallel using $n$ threads). It is instructive to read the paper having in mind as a prime example $s$ being the optimal scheduler (for $n$ threads), which returns a schedule for $n$ threads that minimizes the makespan. However, implementing such a scheduler is computationally prohibitive in practice, so greedy heuristics are typically used instead. To make our results apply even to non-optimal schedulers, we assume a set of minimal, reasonable properties for the scheduler:
  1. (S1)
    Monotonicity in $T$: for any $T\subseteq T^{\prime}$, we have $v^{s}(T)\leq v^{s}(T^{\prime})$. Intuitively, scheduling no fewer transactions takes no less time.
  2. (S2)
    Monotonicity under bundling: consider any set of transactions $T$ and three transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\notin T$, such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$, then we have $v^{s}(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})\leq v^{s}(T\cup\{\textit{tx}_{3}\})$. Intuitively, replacing two transactions by their concatenation makes scheduling no easier.
  3. (S3)
    Monotonicity in $t$ and $K$: consider any set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$, i.e., $\textit{tx}_{1}\lesssim\textit{tx}_{2}$, then we have $v(T\cup\{\textit{tx}_{1}\})\leq v(T\cup\{\textit{tx}_{2}\})$. Intuitively, “larger” transactions are no easier to schedule.
  4. (S4)
    Empty set: $v(\varnothing)=0$.
The optimal scheduler can be easily seen to satisfy these properties. [^11] All but the second straightforward, while for the second, any admissible schedule for $T\cup\{\textit{tx}_{3}\}$ can be turned into a same-makespan admissible schedule for $T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}$ by replacing $\textit{tx}_{3}$ by $\textit{tx}_{1}$ followed immediately by $\textit{tx}_{2}$. Our positive results will apply to any scheduler satisfying the properties, while our negative results will be for the optimal scheduler itself. Henceforth, we drop the $s$ superscript for brevity and follow this convention for who $s$ is.
An attentive reader may notice that (S1)(S3) partially resemble Properties˜1, 2, 3, 4 and 5 for GCMs. This resemblance is not coincidental, but it is important to emphasize that they address different aspects: the former are properties of the scheduler, while the latter are properties of the GCM. A GCM can, in fact, satisfy Properties˜1, 2, 3, 4 and 5 without depending on the scheduler at all (e.g., the current mechanism). Conversely, for mechanisms defined in terms of the scheduler, (S1)(S3) play a crucial role in establishing their properties, including Properties˜1, 2, 3, 4 and 5.
Armed as such, we now return to our latest goal: formalizing the idea that a transaction’s gas consumption should fairly reflect its impact on execution time. We do this as follows:
Property 6 (Scheduling Monotonicity).
Given a set of transactions $T$ and two transactions $\textit{tx}_{1}$ and $\textit{tx}_{2}$, both not in $T$, such that $v(T\cup\{\textit{tx}_{1}\})<v(T\cup\{\textit{tx}_{2}\})$: [^12] Interestingly, unlike our earlier properties, writing $v(T\cup\{\textit{tx}_{1}\})\leq v(T\cup\{\textit{tx}_{2}\})$ here may be too strong, as applying the property twice would then imply that if $v(T\cup\{\textit{tx}_{1}\})=v(T\cup\{\textit{tx}_{2}\})$, then $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})=\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2})$, which is not necessarily desirable. Writing $v(T\cup\{\textit{tx}_{1}\})\leq v(T\cup\{\textit{tx}_{2}\})$ would also unintentionally make the strict and non-strict versions of the property incomparable.
$\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$
Intuitively, transactions with higher marginal contributions to the execution time should consume no less gas. [^13] Technically, what we wrote above in Property 6 are not marginal contributions, but can be made to be by subtracting $v(T)$ from both sides of the inequality in the antecedent. As usual, we define a strict version of this property, where the latter inequality also becomes strict (i.e., higher marginal contributions in execution time imply higher gas consumption). One might be tempted to believe that Scheduling Monotonicity implies Storage Key-Time Monotonicity. However, proving this requires a non-strict inequality in the premise $v(T\cup\{\textit{tx}_{1}\})<v(T\cup\{\textit{tx}_{2}\})$.
There is a second way in which gas consumption should adequately indicate the effort required to execute transactions (Property˜7). Specifically, the gas consumption of all transactions in a block should collectively account for the total time needed to execute the block. This can be seen as properly tracking the time consumed by the execution environment to execute the block.
Property 7 (Efficiency).
Consider a set of transactions $T$ and recall the definition $\textsf{gas}_{T}(T)=\sum_{\textit{tx}\in T}\textsf{gas}_{T}(\textit{tx})$, then:
$\textsf{gas}_{T}(T)=v(T).$
Next, we introduce two practical properties. The first requires that transaction submitters be able to estimate a transaction’s gas consumption in advance. We formalize this by requiring that $\textsf{gas}_{T}(\textit{tx})$ does not depend on $T$ (Property˜8).
Property 8 (Easy Gas Estimation).
Given two sets of transactions $T_{1}$ and $T_{2}$ and a transaction tx belonging to neither $T_{1}$ nor $T_{2}$:
$\textsf{gas}_{T_{1}\cup\{\textit{tx}\}}(\textit{tx})=\textsf{gas}_{T_{2}\cup\{\textit{tx}\}}(\textit{tx}).$
This property ensures a good user experience by making it easy for users to estimate gas usage. If the gas consumption of a transaction depended on the remaining transactions in the block, this estimation would become significantly more complex. In general, we aim to keep gas estimation straightforward to avoid giving an advantage to more sophisticated users. Additionally, as we will see later, a GCM satisfying this property can be seamlessly composed with existing TFMs, retaining their desirable properties (see Section˜6). Sadly, as one might have already guessed, Easy Gas Estimation is incompatible with Scheduling Monotonicity (except for constant mechanisms, i.e., GCMs that return a constant independent of $T$ and tx), and also incompatible with Efficiency:
Theorem 3.4.
Easy Gas Estimation (Property˜8) is:
  1. 1.
    Incompatible with Scheduling Monotonicity (Property˜6) unless using a constant GCM. [^14] Constant GCMs satisfy Scheduling Monotonicity, but not Strict Scheduling Monotonicity.
  2. 2.
    Incompatible with Efficiency (Property˜7).
Proof 3.5.
Consider a mechanism $M$ with Easy Gas Estimation; i.e., $\textsf{gas}^{M}(\textit{tx})$ is meaningful without a subscript for the set of transactions $T$ in the block. Then:
  1. 1.
    Assume that $M$ has Scheduling Monotonicity; we will show that $M$ is a constant mechanism.
    Call two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$ incomparable if neither $K_{1}\subseteq K_{2}$ nor $K_{2}\subseteq K_{1}$. As a first step, we will show that if $\textit{tx}_{1}$ and $\textit{tx}_{2}$ are incomparable, then $\textsf{gas}^{M}(\textit{tx}_{1})=\textsf{gas}^{M}(\textit{tx}_{2})$. To show this, assume that $K_{1}\not\subseteq K_{2}$. We will show that then $\textsf{gas}^{M}(\textit{tx}_{1})\geq\textsf{gas}^{M}(\textit{tx}_{2})$. Exchanging the roles of $\textit{tx}_{1}$ and $\textit{tx}_{2}$ will then give the conclusion. Let $k\in K_{1}\setminus K_{2}$ and $\textit{tx}_{3}$ be another transaction such that $\textit{tx}_{3}\simeq(t_{3},\{k\})$ for some $t_{3}>t_{2}$, and define $T=\{\textit{tx}_{3}\}$. Then, for any number of threads $n\geq 2$ and any scheduler that is optimal for two transactions, we have $v(T\cup\{\textit{tx}_{1}\})=t_{1}+t_{3}>t_{3}=v(T\cup\{\textit{tx}_{2}\})$. By Scheduling Monotonicity, this implies that $\textsf{gas}^{M}(\textit{tx}_{1})\geq\textsf{gas}^{M}(\textit{tx}_{2})$.
    We now know that $M$ associates the same gas consumption to any pair of incomparable transactions. Armed as such, consider two arbitrary transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$. Let $k$ be an arbitrary storage key not in $K_{1}\cup K_{2}$ and $\textit{tx}_{3}$ be a transaction such that $\textit{tx}_{3}\simeq(1,\{r\})$. Since $K_{1}$ and $K_{2}$ are non-empty, one can easily see that $\textit{tx}_{3}$ is incomparable with both $\textit{tx}_{1}$ and $\textit{tx}_{2}$, from which $\textsf{gas}^{M}(\textit{tx}_{1})=\textsf{gas}^{M}(\textit{tx}_{3})=\textsf{gas}^{M}(\textit{tx}_{2})$.
  2. 2.
    Assume for a contradiction that $M$ has Efficiency, then for any transaction $\textit{tx}\simeq(t,K)$ we have $\textsf{gas}^{M}(\textit{tx})=\textsf{gas}_{\varnothing}^{M}(\{\textit{tx}\})=v(\{\textit{tx}\})=t$. Hence, by definition, $M$ is the current mechanism, which can be easily seen to not satisfy Efficiency: assume $n\geq 2$ threads and consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2}\}$ where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$ and $\textit{tx}_{2}\simeq(1,\{k_{2}\})$. In this case, $\textsf{gas}^{M}(T)=\textsf{gas}^{M}(\textit{tx}_{1})+\textsf{gas}^{M}(\textit{tx}_{1})=1+1=2\neq 1=v(T)$.
Our second practical property requires that gas consumption be efficiently computable, i.e., in polynomial time (Property˜9). A GCM that does not satisfy this property would be unsuitable for the blockchain context, where gas computation is intended to be a straightforward component.
Property 9 (Poly-time Computable).
There exists a polynomial-time algorithm that takes as input a transaction set $T$ and a transaction tx and outputs $\textsf{gas}_{T\cup\{\textit{tx}\}}(\textit{tx})$.
4 GCM Proposals
We now propose multiple GCM designs. Motivated by the incompatibilities in Theorem˜3.4, these fall into two categories:
  1. (C1)
    Mechanisms with Easy Gas Estimation, in which each transaction’s gas consumption is computed in isolation. Such mechanisms tend to be both straightforward and attractive but can achieve neither Scheduling Monotonicity [^15] Except for constant GCMs. nor Efficiency.
  2. (C2)
    Mechanisms without Easy Gas Estimation, which, given a set of transactions $T$, rely on $v(T)$ or, more generally, $\left(v(T^{\prime})\right)_{T^{\prime}\subseteq T^{\prime}}$ to holistically calculate gas consumptions for the block $T$, requiring knowledge of the entire block. Instead, these mechanisms aim to achieve Scheduling Monotonicity and/or Efficiency.
4.1 Mechanisms with Easy Gas Estimation
We already introduced the current GCM (Definition˜2.3). For the next mechanism, we assume a globally available vector of positive weights for the storage keys $(w_{k})_{k\in\mathcal{K}}$. For instance, these weights could all be 1. Alternatively, higher-weight storage keys could correspond to storage keys under higher (historical) demand. For our purposes, we only need to assume that the weights are known and do not depend on the block being built. In practice, one could update the weights between blocks to accurately reflect storage key demand (the exact details are not relevant here; see Section˜6 for further discussion). Given the weights, we define:
Definition 4.1 (Weighted Area GCM).
Given a set of transactions $T$ and a transaction $\textit{tx}\in T$ with $\textit{tx}\simeq(t,K)$, the Weighted Area GCM computes the amount of gas used by tx as follows:
$\textsf{gas}_{T}^{\textit{WA}}(\textit{tx}):=t\cdot\left(1+\sum_{k\in K}w_{k}\right)$ (1)
Since this does not depend on $T$, we often drop the subscript.
To gain intuition, it is instructive to consider the former case, where all weights are 1, i.e., the unweighted area mechanism, in which case Equation˜1 becomes $\textsf{gas}_{T}^{\textit{WA}}(\textit{tx})=t\cdot\left(1+|K|\right)$. This mechanism can also be viewed as the addition of two terms: the current term, i.e., the gas consumption of tx under the current GCM, namely $t$, and the area term, i.e., the area of a $t\times|K|$ rectangle (which is also how we draw transactions in our diagrams). The current term is helpful to ensure no transaction ever costs too little when the weights of all storage keys accessed by a transaction are too close to zero. [^16] This term has a practical motivation. In reality, transactions will be scheduled on a small finite number of threads. Thus, even if they do not access any high-demand storage key, they occupy space in the schedule proportional to their execution time. Note, moreover, that this term could technically be omitted, and our results would still hold.
[…]
4.2 Mechanisms without Easy Gas Estimation
We now switch gears towards mechanisms without Easy Gas Estimation that instead aim to achieve Scheduling Monotonicity and/or Efficiency. Intuitively, for a set of transactions $T$, such mechanisms should start from $v(T)$ and take each transaction’s marginal contribution towards $v(T)$ as its gas consumption. This, however, has to be done carefully, as, e.g., simply looking for each transaction $\textit{tx}\in T$ at $v(T)-v(T\setminus\{\textit{tx}\})$ does not help. Note that one can easily construct cases where removing any one transaction from $T$ does not change $v(T)$, so all reported numbers will be 0. However, all these numbers being zero does not mean that no transaction contributes to $v(T)$; it just means that we also have to consider removing multiple transactions to see the differences. This gives us the idea to look at the marginal contribution of each transaction $\textit{tx}\in T$ when added to each possible subset $S\subseteq T\setminus\{\textit{tx}\}$, and not just to $S=T\setminus\{\textit{tx}\}$ as before. Formally, we would like to compute tx’s gas consumption as an aggregate of $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$. We next present two mechanisms based on this idea: the Shapley (Definition˜4.2) and Banzhaf (Definition˜4.4) GCMs, inspired by corresponding concepts in cooperative game theory.
Definition 4.2 (Shapley GCM).
Given a set of transactions $T$ consisting of $|T|=n$ transactions and a transaction $\textit{tx}\in T$, the Shapley GCM computes the amount of gas used by tx as follows:
$\textsf{gas}_{T}^{S}(\textit{tx}):=$ $\frac{1}{n!}\sum_{\sigma}\left[v(P_{\textit{tx}}^{\sigma}\cup\{\textit{tx}\})-v(P_{\textit{tx}}^{\sigma})\right]$ (2)
$=$ $\sum_{S\subseteq T\setminus\{\textit{tx}\}}\frac{|S|!\cdot(n-|S|-1)!}{n!}\left[v(S\cup\{\textit{tx}\})-v(S)\right].$ (3)
Here, $\sigma$ ranges over the $n!$ possible ways to order the transactions in $T$ and $P_{\textit{tx}}^{\sigma}$ denotes the set of transactions that precede tx in the order $\sigma$. The equality between Equations˜2 and 3 follows by counting the number of orders $\sigma$ such that $P_{\textit{tx}}^{\sigma}=S$, which there are $|S|!\cdot(n-|S|-1)!$ of.
Another way to understand the Shapley GCM is through the following probabilistic experiment:
  1. 1.
    Select an ordering $\sigma$ of the transactions in $T$ uniformly at random.
  2. 2.
    Start with an empty set of transactions and add transactions one by one in the order given by $\sigma$.
  3. 3.
    Whenever a transaction tx is added, let $S=P_{\textit{tx}}^{\sigma}$ be the set of transactions just before adding it and compute tx’s marginal contribution to the execution time as $v(S\cup\{tx\})-v(S)$; i.e., by how much did the execution time increase by adding tx to the current set of transactions.
  4. 4.
    The gas consumption of $\textit{tx}\in T$ is the expectation of its marginal contribution across the experiment.
The reader familiar with cooperative game theory will have already recognized the immediate connection with Shapley values: if we see each transaction $\textit{tx}\in T$ as a player in a game with valuation function $v:2^{T}\to\mathbb{R}$, then $\textsf{gas}_{T}^{S}(\textit{tx})$ is precisely the celebrated Shapley value of player tx, more traditionally written $\phi_{\textit{tx}}(v)$. One classical property of Shapley values is that, given $v(\varnothing)=0$, as is the case for us by (S4), their sum equals the valuation of the grand coalition $T$: $\sum_{\textit{tx}\in T}\phi_{\textit{tx}}(v)=v(T)$. The proof is straightforward: for any fixed ordering $\sigma$ of $T$, the sum of the marginal contributions of the transactions is a telescoping sum, i.e., except first and last terms, all others appear once positively and once negatively, canceling as a result and leaving us with $v(T)-v(\varnothing)=v(T)$. Since the sum does not depend on $\sigma$, the same holds when taking the expectation with respect to $\sigma$. This already gives us our first property of the Shapley GCM, namely Efficiency:
Lemma 4.3.
The Shapley GCM satisfies Efficiency (Property˜7).
We postpone studying the Shapley GCM further until introducing our other mechanisms in the section.
A second GCM based on the idea that a transaction’s gas consumption should be its marginal contribution to the execution time is the Banzhaf mechanism:
Definition 4.4 (Banzhaf GCM).
Given a set of transactions $T$ consisting of $|T|=n$ transactions and a transaction $\textit{tx}\in T$, the Banzhaf GCM computes the amount of gas used by tx as follows:
$\textsf{gas}_{T}^{B}(\textit{tx}):=\frac{1}{2^{n-1}}\sum_{S\subseteq T\setminus\{\textit{tx}\}}\left[v(S\cup\{\textit{tx}\})-v(S)\right].$
As for the Shapley mechanism, the Banzhaf mechanism can be understood probabilistically, but this time with a separate experiment for each transaction tx. Sample a subset $S$ of transactions other than tx uniformly at random and compute tx’s marginal contribution to the execution time when added to $S$, namely $v(S\cup\{\textit{tx}\})-v(S)$. The gas consumption of tx is then its expected marginal contribution to the execution time. The familiar reader will recognize this as the definition of the Banzhaf power index $\beta_{\textit{tx}}(v)$. For consistency with Shapley values, we will instead call these Banzhaf values. While more straightforward than the corresponding experiment used in defining Shapley values, the fact that we now have $n$ separate experiments makes the sum of the values no longer well-behaved, losing Efficiency for the Banzhaf mechanism:
Lemma 4.5.
The Banzhaf GCM does not satisfy Efficiency (Property˜7).
We conclude this section by introducing two additional reasonable mechanisms without Easy Gas Estimation. These mechanisms are notably more straightforward than the Shapley and Banzhaf GCMs, as they avoid computing marginal contributions. However, they have other drawbacks that will become more apparent when we begin studying their normative properties alongside the other mechanisms.
Definition 4.6 (Time-Proportional Makespan GCM).
Given a set of transactions $T$ and a transaction $\textit{tx}\in T$, the Time-Proportional Makespan (TPM) GCM computes the amount of gas used by tx as follows:
$\textsf{gas}_{T}^{\textit{TPM}}(\textit{tx}):=\frac{t(\textit{tx})}{\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})}\cdot v(T).$
Definition 4.7 (Equally-Split Makespan GCM).
Given a set of transactions $T$ and a transaction $\textit{tx}\in T$, the Equally-Split Makespan (ESM) GCM computes the amount of gas used by tx as follows:
$\textsf{gas}_{T}^{\textit{ESM}}(\textit{tx}):=\frac{v(T)}{|T|}.$
We also consider the following rather pathological mechanism because of the combination of properties it turns out to satisfy (that none of our other mechanisms do):
Definition 4.8 (Exponentially-Split Makespan GCM).
Given a set of transactions $T$ and a transaction $\textit{tx}\in T$, the Exponentially-Split Makespan (XSM) GCM computes the amount of gas used by tx as follows:
$\textsf{gas}_{T}^{\textit{XSM}}(\textit{tx}):=\frac{v(T)}{3^{|T|}}.$
5 Analysis of Our GCMs
In this section, we provide a detailed analysis of the normative properties of our GCMs. Our results are summarized in Table˜1.
Property Current W. Area Shapley Banzhaf TPM ESM XSM
Storage Key Monotonicity (Property˜1) == << $\leq$ $\leq$ $\leq$ $\leq$ $\leq$
Time Monotonicity (Property˜2) << << << << << $\leq$ $\leq$
Storage Key-Time Monotonicity (Property˜3) $\leq$ << $\leq$ $\leq$ $\leq$ $\leq$ $\leq$
[…]
Transaction Bundling (Property˜5) == $\leq$ $\leq$ $\leq$ <<
Scheduling Monotonicity (Property˜6) << <<
Efficiency (Property˜7)
Easy Gas Estimation (Property˜8)
Poly-time Computable (Property˜9) $S(v)$ $B(v)$ $v$ $v$ $v$
Table 1: Comparison of GCMs based on their adherence to the defined properties. <\,<\, indicates that the mechanism strictly satisfies the property, =\,=\, indicates trivial satisfaction (by equality), and $\leq\,$ indicates satisfaction (not necessarily strict). A $\checkmark$” means the property is satisfied; $\times$” means it is not satisfied. For computational complexity, “vv” means the mechanism is as hard to compute as $v(\cdot)$ itself, “S(v)S(v)” means it is as hard as computing Shapley values for $v$, and “B(v)B(v)” means it is as hard as computing Banzhaf values for $v$.
5.1 Mechanisms with Easy Gas Estimation
In this section, we analyze the current and Weighted Area GCMs. By definition, both satisfy Easy Gas Estimation (Property˜8) and are Poly-time Computable (Property˜9). Since they satisfy Easy Gas Estimation but are not constant GCMs, Theorem˜3.4 implies that neither satisfies Scheduling Monotonicity (Property˜6) nor Efficiency (Property˜7). Furthermore, both mechanisms satisfy strict Time Monotonicity (Property˜2): increasing the time $t$ of a transaction strictly increases its gas consumption. Similarly, both satisfy strict Set Inclusion (Property˜4): since transactions’ gas consumptions are strictly positive, a strict superset of a given set of transactions consumes strictly more gas. For the remaining three properties, the two mechanisms behave slightly differently:
Storage Key Monotonicity (Property˜1). The current mechanism ignores the set of storage keys $K$ that a transaction accesses, so replacing $K$ with a strict superset of it does not change the transaction’s gas consumption. Therefore, the current mechanism satisfies Storage Key Monotonicity with equality. On the other hand, the Weighted Area mechanism adds an extra term of $t\cdot w_{k}>0$ to the gas consumption for each additional storage key $k$ added to $K$, so it satisfies strict Storage Key Monotonicity.
Storage Key-Time Monotonicity (Property˜3). From the above results on whether our two mechanisms satisfy Storage Key Monotonicity and Time Monotonicity, Lemmas˜3.1 and 3.2 allow us to conclude that the current mechanism satisfies Storage Key-Time Monotonicity, while the Weighted Area mechanism satisfies it strictly.
Transaction Bundling (Property˜5). Concatenating two transactions with times $t_{1}$ and $t_{2}$ results in a transaction with time $t_{1}+t_{2}$. Since the current mechanism equates gas consumption with time, the bundled transaction has the same gas consumption as the two individual transactions combined. This implies that the current mechanism satisfies Transaction Bundling with equality. Finally, the Weighted Area mechanism satisfies Transaction Bundling, as shown below. If we restrict ourselves to bundling transactions with different storage key sets, the property is strictly satisfied.
Lemma 5.1.
The Weighted Area GCM satisfies Transaction Bundling (Property˜5). If we only consider bundling transactions with different storage key sets, the property is strictly satisfied.
Proof 5.2.
Consider two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$. Let $\textit{tx}_{3}\simeq(t_{1}+t_{2},K_{1}\cup K_{2})$ be a transaction consisting of the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$. We want to show that $\textsf{gas}^{\textit{WA}}(\textit{tx}_{1})+\textsf{gas}^{\textit{WA}}(\textit{tx}_{2})\leq\textsf{gas}^{\textit{WA}}(\textit{tx}_{3})$. By definition, this amounts to:
$t_{1}\cdot\left(1+\sum_{k\in K_{1}}w_{k}\right)+t_{2}\cdot\left(1+\sum_{k\in K_{2}}w_{k}\right)\leq(t_{1}+t_{2})\cdot\left(1+\sum_{k\in K_{1}\cup K_{2}}w_{k}\right)$ (4)
Which is true because $\sum_{k\in K_{i}}w_{k}\leq\sum_{k\in K_{1}\cup K_{2}}w_{k}$ for $i\in\{1,2\}$. Because the weights are strictly positive, equality occurs if and only if $K_{1}=K_{1}\cup K_{2}$ and $K_{2}=K_{1}\cup K_{2}$; i.e., $K_{1}=K_{2}$. Hence, the property is satisfied strictly if we restrict bundling transactions to cases where $K_{1}\neq K_{2}$.
5.2 Mechanisms without Easy Gas Estimation
In this section, we analyze the Shapley, Banzhaf, TPM, ESM, and XSM GCMs. By definition, all of them require knowledge of $T$ to compute $\textsf{gas}_{T}(\textit{tx})$, so they do not satisfy Easy Gas Estimation (Property˜8). Next, we examine each of the remaining properties individually and analyze whether they hold for our five mechanisms.
Poly-time Computable (Property˜9). The TPM, ESM, and XSM GCMs are all Poly-time Computable whenever determining the makespan of a given set of transactions under the chosen scheduler is feasible in polynomial time, i.e., when $v(T)$ can be computed in polynomial time. [^17] Notably, this is essentially never true for non-trivial makespan minimization problems. In particular, if we restrict our attention to the case of unit-length transactions (i.e., with $t=1$) and infinitely many threads $n=\infty$, then checking for the existence of a schedule with makespan $c$ corresponds to checking whether the intersection graph of the transactions (i.e., where edges correspond to transactions with intersecting storage key sets) is $c$-colorable. For $c=3$, this is well-known to be NP-complete, but this result is for general graphs. However, there is a straightforward way to model any graph $G=(V,E)$ as an intersection graph of transactions: vertices $v\in V$ are transactions and for every edge $(u,v)\in E$, create a new storage key $k$ and add it to the storage key sets of transactions $u$ and $v$. In fact, the problems are all equally difficult to computing such makespans. In contrast, computing gas consumptions under the Shapley and Banzhaf GCMs hits another hurdle: it is no longer enough to be able to efficiently compute $v(T)$, but instead, one needs to be able to compute the Shapley or Banzhaf values of the transactions, involving aggregating over $(v(T^{\prime}))_{T^{\prime}\subseteq T}$. Hence, computing Shapley and Banzhaf values is, in general, NP-hard and #P-complete [18, 23, 40, 41, 51]. Moreover, no deterministic polynomial-time algorithm can approximate the Shapley values within a constant factor unless P = NP. Using randomization could, in principle, circumvent this: to approximate an average consisting of exponentially many terms, sample polynomially many uniformly at random, and take their average. One could imagine implementing this in a blockchain via a VRF or similar mechanism to ensure unbiased randomness. Still, this comes with a notable downside: the accuracy of the computed gas consumption may be hard to predict in advance and might vary wildly. Moreover, even with randomization, one could still run into issues computing the sampled terms, which is as hard as evaluating $v$ a constant number of times. See [14] for an ampler discussion of computing Shapley and Banzhaf values. Note that the previous results mostly pertain to functions $v$ of a different shape than ours (i.e., not related to scheduling). We have not attempted to show that hardness is retained in our context, but expect this to be true. Note still that because the Shapley GCM is Efficient, it can be used to compute $v(T)$ by adding up the gas consumptions of all transactions in $T$, meaning that the Shapley GCM is at least as difficult to compute as $v(T)$. However, this simple reduction no longer works for the Banzhaf GCM.
Efficiency (Property˜7). We have already seen that the Shapley GCM is Efficient (Lemma˜4.3) while the Banzhaf GCM is not (Lemma˜4.5). Moreover, by adding up the gas consumption, one can immediately see from the definitions that TPM and ESM are Efficient, while XSM is not. We take this occasion to also note that any GCM can be made Efficient by appropriately scaling its output. Most notably, applying this to the Banzhaf GCM yields a mechanism based on the well-known normalized Banzhaf values, though the resulting mechanism otherwise seems rather poorly behaved.
Scheduling Monotonicity (Property˜6). The ESM and XSM GCMs can be easily seen to strictly satisfy Scheduling Monotonicity. This is because, under both mechanisms, given a set of transactions $T$, the gas consumption of any transaction in $T$ is computed as $f(|T|)\cdot v(T)$, where $f$ is either $\frac{1}{x}$ or $\frac{1}{3^{x}}$. Notably, the first factor depends only on $|T|$, so if a transaction in $T$ were modified in a way that increases the makespan, this increase would also be reflected proportionally in its gas consumption. In contrast, the Shapley, Banzhaf, and TPM GCMs do not satisfy Scheduling Monotonicity (proven in Lemma˜5.3 below). This may be particularly surprising for the Shapley and Banzhaf GCMs, as they were specifically designed to account for marginal increases in makespan. The catch is that Scheduling Monotonicity considers replacing a transaction tx with another one leading to an increase in makespan. However, some elements in $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$ may still decrease as a result (except for $S=T\setminus\{\textit{tx}\}$, for which we assumed an increase). Because both the Shapley and Banzhaf values of tx take a weighted average over these values, the average might still decrease, which is what happens in the example in the lemma below.
Lemma 5.3.
The Shapley, Banzhaf and TPM GCMs do not satisfy Scheduling Monotonicity (Property˜6).
Proof 5.4.
Consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2}\}$ and two additional transactions $\textit{tx}_{3}$ and $\textit{tx}_{4}$ such that $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(3,\{k_{2}\})$, $\textit{tx}_{3}\simeq(2,\{k_{1}\})$, and $\textit{tx}_{4}\simeq(1,\{k_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler we have $v(T\cup\{\textit{tx}_{3}\})=3<4=v(T\cup\{\textit{tx}_{4}\})$. However:
$\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{S}(\textit{tx}_{3})=\frac{2+2+2+0+0+0}{6}=1>$ $\frac{5}{6}=\frac{1+1+1+1+1+0}{6}=\textsf{gas}_{T\cup\{\textit{tx}_{4}\}}^{S}(\textit{tx}_{4})$
$\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{B}(\textit{tx}_{3})=\frac{2+2+0+0}{4}=1>$ $\frac{3}{4}=\frac{1+0+1+1}{4}=\textsf{gas}_{T\cup\{\textit{tx}_{4}\}}^{B}(\textit{tx}_{4})$
$\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{\textit{TPM}}(\textit{tx}_{3})=\frac{2}{1+3+2}\cdot 3=1>$ $\frac{4}{5}=\frac{1}{1+3+1}\cdot 4=\textsf{gas}_{T\cup\{\textit{tx}_{4}\}}^{\textit{TPM}}(\textit{tx}_{4}).$
So, the Shapley, Banzhaf, and TPM GCMs all violate Scheduling Monotonicity in this case.
Transaction Bundling (Property˜5). We find that the Banzhaf, TPM and XSM GCMs satisfy Transaction Bundling, only XSM satisfying it strictly, while the Shapley and ESM GCMs do not satisfy the property. We prove these facts in the following 5 lemmas.
Lemma 5.5.
The Shapley GCM does not satisfy Transaction Bundling (Property˜5).
Lemma 5.6.
The Banzhaf GCM satisfies Transaction Bundling (Property˜5).
Lemma 5.7.
The TPM GCM satisfies Transaction Bundling (Property˜5).
Lemma 5.8.
The ESM GCM does not satisfy Transaction Bundling (Property˜5).
Lemma 5.9.
The XSM GCM strictly satisfies Transaction Bundling (Property˜5).
Set Inclusion (Property˜4). We find that the TPM and ESM GCMs satisfy Set Inclusion, while the Shapley, Banzhaf and XSM GCMs do not satisfy the property. We prove these facts in the following 5 lemmas.
Lemma 5.10.
The Shapley GCM does not satisfy Set Inclusion (Property˜4).
Lemma 5.11.
The Banzhaf GCM does not satisfy Set Inclusion (Property˜4).
Lemma 5.12.
The TPM GCM satisfies Set Inclusion (Property˜4).
Lemma 5.13.
The ESM GCM satisfies Set Inclusion (Property˜4).
Lemma 5.14.
The XSM GCM does not satisfy Set Inclusion (Property˜4).
Storage Key-Time Monotonicity (Property˜3). All five mechanisms satisfy this property. For the ESM and XSM GCMs, this is an immediate consequence of property (S3). For the Shapley and Banzhaf GCMs, one can see this by recalling that they compute the gas consumption of a transaction $\textit{tx}\in T$ as a weighted average over $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$. Hence, by property (S3), when tx is replaced with some $\textit{tx}^{\prime}\gtrsim\textit{tx}$, no term in the previous decreases, so their weighted average also does not decrease. Finally, for the TPM GCM, we show this in the lemma after the next paragraph*.
Storage Key Monotonicity (Property˜1) and Time Monotonicity (Property˜2). Because all five mechanisms satisfy Storage Key-Time Monotonicity, by Lemma˜3.1, they also all satisfy Storage Key Monotonicity and Time Monotonicity. Out of the five mechanisms, the Shapley, Banzhaf, and TPM GCMs satisfy the property strictly. For the first two, this is because when $t(\textit{tx})$ increases, at least one term in $\left(v(S\cup\{\textit{tx}\})-v(S)\right)_{S\subseteq T\setminus\{\textit{tx}\}}$ strictly increases, namely the term for $S=\varnothing$, while no terms decrease by property (S3). Last, for the TPM GCM, we show this in the lemma below.
Lemma 5.15.
The TPM GCM satisfies Storage Key-Time Monotonicity (Property˜3) and strict Time Monotonicity (Property˜2).
6 Towards a Fee Market for Parallel Execution
Armed with an understanding of the trade-offs between desirable properties in a GCM, in particular the impossibility of satisfying all properties within a single mechanism, and informed by our analysis of various candidate mechanisms and their properties, we propose two mechanisms for practical implementations of a fee market supporting parallel execution. Each represents one side of the design spectrum: one drawn from the class of mechanisms with Easy Gas Estimation, and the other from the class without it.
The advantage of adopting a mechanism with Easy Gas Estimation is that satisfying this property ensures that each transaction consumes a fixed amount of gas, regardless of other transactions in the same block. From the perspective of currently deployed TFMs (e.g., EIP-1559), which process transactions with fixed sizes, nothing changes. Thus, the existing properties of these mechanisms remain intact. On a high level, the key properties we strive for in a TFM are incentive compatibility for both block producers and users, welfare optimality, and collusion resistance. However, no TFM can achieve all these properties simultaneously [17, 16, 26]. Importantly, when composing a GCM that satisfies Easy Gas Estimation with a TFM of choice, the level of sophistication required from users in their bidding strategy does not increase, unlike in currently deployed fee markets for parallel execution [39, 43].
On the other hand, foregoing Easy Gas Estimation makes it possible to price transactions according to the load they impose on the network relative to the other transactions in the block, rather than evaluating each one in isolation. Mechanisms in this class can align the total gas charged in a block with the actual execution cost of that block, enabling resource pricing at the block level rather than only at the transaction level. Recent discussions in Ethereum research have explored similar designs under the umbrella of block-level fee markets [10]. If the protocol is already moving toward block-level metering for other resources, such as data availability or storage contention, extending this approach to execution costs could enable direct efficiency gains from parallel execution that are not achievable with transaction-level pricing alone.
For the class of mechanisms with Easy Gas Estimation, the natural choice is the Weighted Area GCM. Given that it satisfies Easy Gas Estimation, it achieves the most additional properties one could hope for in a (non-constant) GCM (see Theorem˜3.4).
For the class of mechanisms without Easy Gas Estimation, we propose the Time-Proportional Makespan GCM as a concrete candidate. By allocating gas in proportion to each transaction’s execution time relative to the block’s total makespan, it provides a simple way to align gas consumption with how transactions constrain parallel execution within a block.
The decision between these approaches ultimately depends on protocol-level priorities: maintaining user-side simplicity and compatibility with existing TFMs through Easy Gas Estimation, which also makes implementation simpler by reusing current transaction-level pricing infrastructure, or pursuing block-level efficiency and more accurate resource pricing at the cost of giving up the previous benefits.
7 Related Work
7.1 Transaction Fee Mechanisms
There is extensive research on blockchain fee markets, with a particular focus on Ethereum and Bitcoin. Early studies primarily examined Bitcoin, exploring monopolistic pricing mechanisms [36, 63]. More recent contributions to this field include [47, 25, 49]. Unlike these works, our study concerns measuring storage key usage on a blockchain with client-side parallel execution, rather than focusing on pricing.
The TFM design framework was introduced by Roughgarden [53, 54]. Roughgarden’s analysis of the EIP-1559 mechanism [11] initiated an active line of research on TFMs. Chung and Shi [17] demonstrated that no TFM can be ideal — meaning it cannot simultaneously be incentive-compatible for users and block producers while also being resistant to collusion between the two. This conclusion holds even for weaker definitions of collusion resilience, as shown by Chung et al. [16] and Gafni and Yaish [26]. Finally, attempts to address these limitations using cryptographic techniques [58, 61] have made progress in overcoming certain impossibilities, while other attempts relax the desiderata [27]. However, designing an ideal TFM still remains out of reach. While these studies examine the limitations of TFMs, our focus is on GCMs for parallel execution and how to integrate them with a TFM.
A related body of work examines the dynamics of TFMs over multiple blocks, particularly focusing on the base fee in EIP-1559. Leonardos et al. [37, 38] demonstrate that the stability of the base fee depends on the adjustment parameter, with short-term volatility but long-term block size stability. Reijsbergen et al. [52] suggest using an adaptive adjustment parameter to mitigate block size fluctuations, while Ferreira et al. [24] highlight user experience issues caused by bounded base fee oscillations. Additionally, Hougaard and Pourpouneh [31] and Azouvi et al. [5] reveal that the base fee can be manipulated by non-myopic miners.
Given the discussion surrounding multi-dimensional fees in Ethereum [8, 9] and the deployment of EIP-4844 [12] (a first step towards a multi-dimensional fee market on Ethereum), a recent line of work explores multi-dimensional fee markets, focusing on efficient pricing mechanisms and their optimality. This work is further refined by Diamandis et al. [20], who design and analyze multi-dimensional blockchain fee markets to align incentives and improve network performance. Building on this, Angeris et al. [2] prove that such fee markets are nearly optimal, with efficiency improving over time even under adversarial conditions. Multidimensional fee markets are closely related to fee markets designed for parallel execution. In particular, in the weighted area GCM, the weights can be interpreted as fees within a multidimensional fee market. Unlike previous literature on multidimensional fee markets, we focus on parallelization, introduce desirable properties, and evaluate how various mechanisms perform.
Further extensions of TFMs have emerged. Bahrani et al. [7] consider TFMs in the presence of maximal extractable value (MEV), i.e., value extractable by the block producer. Further, Wang et al. [60] design a fee mechanism for proof networks, whereas Bahrani et al. [6] introduce a transaction fee mechanism for heterogeneous computation. Our work most closely relates to the latter, but, in contrast, our chosen approach is closer to multidimensional fee markets, trading complexity for the block producer for stronger incentive compatibility for the user.
Local fee markets have recently been a topic of discussion in the blockchain space [22, 19, 39, 33]. The core idea is that transactions interacting with highly contested states incur higher fees, while those involving non-contested states pay lower fees. However, discussions on local fee markets have largely remained high-level, without a precise characterization of the desired properties beyond this general goal. Moreover, currently implemented local fee markets [39] require significant user sophistication to set fees appropriately. In this work, we formalize the desiderata for fee markets in the context of parallel execution and identify the weighted area GCM as a promising candidate. One key advantage is its compatibility with a TFM, enabling simple fee estimation for users.
7.2 Parallel Execution
Blockchain concurrency has been a focal point in an active line of research. In particular, numerous efforts have aimed to enable parallel transaction processing through speculative execution [57, 64, 1, 21, 3, 28, 15, 55]. Note that speculative execution is already deployed by multiple blockchains [4, 56, 42]. Static analysis has also been employed to identify parallelizable transactions, though it cannot completely eliminate inherent dependencies [50, 44]. Similarly, Neiheiser et al. [46] demonstrate how parallel execution can assist struggling nodes in catching up. While these works are orthogonal to ours, they highlight the overhead of parallel execution when there is no advance knowledge about a transaction’s state accesses.
Further, Saraph and Herlihy [55] and Heimbach et al. [29] have evaluated the parallelization potential of the Ethereum workload. The latter demonstrates that a speedup of approximately fivefold is achievable, assuming state accesses are known in advance. Additionally, Solana [59] and Sui [45] already perform parallel execution with advance knowledge of state accesses. However, in practice, state accesses are not known beforehand on many blockchains such as Ethereum. There, less than 2% of transactions disclose them proactively, as shown by Heimbach et al. [30], due to a lack of incentives. In this work, we aim to take a step toward unlocking the parallelization potential by designing a TFM that supports parallel execution. This mechanism relies on the disclosure of state accesses as done in Solana [59] and Sui [45].
7.3 Cloud Computing Pricing and Parallel Resource Allocation
While our primary focus is on transaction fee mechanisms for blockchain networks, parallels can be drawn to resource allocation and pricing in cloud computing. Several papers study pricing schemes for cloud computing resources in a monopolistic setting, where a single provider sells access to multiple resources and sets prices to optimize revenue [62]. In these models, the provider posts prices for resource usage, users select resource bundles to maximize their individual surplus, and the provider earns revenue from the resulting allocation. Mechanisms such as CloudPack [32] extend this approach by incorporating workload flexibility, using concepts like Shapley values for fair cost distribution among colocated jobs. Similarly, works such as [48] address resource contention through dynamic pricing strategies that adjust costs based on runtime slowdowns.
While structurally similar, the blockchain context introduces an additional constraint. In addition to objectives like maximizing provider revenue, it emphasizes maximizing throughput by ensuring transactions are processed efficiently, minimizing their collective execution time within a block. Moreover, blockchain transactions explicitly declare access to distinct storage keys, introducing contention through overlapping access patterns. This adds complexity compared to cloud models, which typically abstract resource demands as scalar quantities and address contention at a more aggregated level.
Capturing the effects of storage key-level conflicts on parallel execution in blockchains requires fee computation mechanisms that are sensitive to these fine-grained interactions. Thus, while conceptually addressing similar challenges, the blockchain setting imposes unique constraints that motivate the need for specialized GCMs as explored in this work.
8 Conclusion
In this work, we took a step towards creating a fee market that meets the demands of parallel execution environments while also upholding the properties we want from a TFM.
Recently, the idea of local fee markets has been proposed for blockchains that support parallel execution. However, to the best of our knowledge, before this work, the demands on these fee markets have only been outlined at a very high level, and the markets that have been implemented are not ideal yet, e.g., they require high levels of sophistication from users when bidding.
In this work, we addressed this gap by introducing a framework with two key components: a GCM, which measures the execution-related load a transaction imposes on the network in units of gas, and a TFM, which determines the cost associated with each unit of gas. We then formalized the desired properties for the GCM in such a fee market. After outlining these desiderata, we evaluated various mechanisms against them and identified two strong candidates through this analysis: the weighted area GCM for the class of mechanisms with Easy Gas Estimation, and the time-proportional makespan GCM for the class without it.
Setting the right incentives in fee markets for parallel execution is crucial to unlocking the full potential of execution layer parallelization, and we hope that our work contributes to the development of fee markets capable of meeting the demands of such environments.
Appendix A Proofs Omitted From Section˜3
See 3.1
Proof A.1.
Properties˜1 and 2 are special cases of Property˜3, corresponding to the scenarios where $t_{1}=t_{2}$ and $K_{1}=K_{2}$, respectively. Therefore, the $(\Rightarrow)$ direction holds. To prove the $(\Leftarrow)$ direction, assume Properties˜1 and 2 hold and consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$. Let $\textit{tx}_{3}$ be another transaction not in $T$ such that $\textit{tx}_{3}\simeq(t_{1},K_{2})$. By first applying Property˜1 and then Property˜2, we obtain:
$\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3})\leq\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$
This establishes the conclusion.
See 3.2
Proof A.2.
The strict variants of Properties˜1 and 2 are special cases of the strict variant of Property˜3, corresponding to the scenarios where $t_{1}=t_{2}$ and $K_{1}=K_{2}$, respectively. Therefore, the $(\Rightarrow)$ direction holds. To prove the $(\Leftarrow)$ direction, assume Properties˜1 and 2 hold strictly and consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$, and additionally $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$ (i.e., at least one of the previous holds strictly). Let $\textit{tx}_{3}$ be another transaction not in $T$ such that $\textit{tx}_{3}\simeq(t_{1},K_{2})$. By first applying Property˜1 and then Property˜2 (their non-strict versions), we obtain:
$\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3})\leq\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$ (5)
From this, we get that $\textsf{gas}_{T\cup{\textit{tx}_{1}}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup{\textit{tx}_{2}}}(\textit{tx}_{2})$, so it remains to rule out the equality case. Assume for a contradiction that $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})=\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2})$, from which the two inequalities hold with equality in Equation˜5. Because Properties˜1 and 2 hold strictly, this can only be the case if $\textit{tx}_{1}\simeq\textit{tx}_{3}\simeq\textit{tx}_{2}$, contradicting the assumption that $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$.
Appendix B Proofs Omitted From Section˜4
See 4.5
Proof B.1.
Consider the set of transactions $T=\{\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\}$ where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(1,\{k_{1}\})$, and $\textit{tx}_{3}\simeq(1,\{k_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
$v(\{\textit{tx}_{1}\})=v(\{\textit{tx}_{2}\})=v(\{\textit{tx}_{3}\})=v(\{\textit{tx}_{1},\textit{tx}_{3}\})=v(\{\textit{tx}_{2},\textit{tx}_{3}\})=1,\quad v(\{\textit{tx}_{1},\textit{tx}_{2}\})=v(T)=2.$
A short calculation then shows that:
$\textsf{gas}_{T}^{B}(\textit{tx}_{1})=\textsf{gas}_{T}^{B}(\textit{tx}_{2})=\frac{1+1+0+1}{4}=\frac{3}{4},\quad\textsf{gas}_{T}^{B}(\textit{tx}_{3})=\frac{1+0+0+0}{4}=\frac{1}{4}.$
Thus, $\sum_{\textit{tx}\in T}\textsf{gas}_{T}^{B}(\textit{tx})=\frac{3}{4}+\frac{3}{4}+\frac{1}{4}=\frac{7}{4}\neq 2=v(T)$, violating Efficiency.
Appendix C Proofs Omitted From Section˜5
C.1 Transaction Bundling (Property˜5)
See 5.5
Proof C.1.
Consider the set of transactions $T=\{\textit{tx}_{4}\}$ and three other transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$, where $\textit{tx}_{1}\simeq(1,\{k_{2}\})$, $\textit{tx}_{2}\simeq(1,\{k_{2}\})$, $\textit{tx}_{3}\simeq(2,\{k_{2}\})$, and $\textit{tx}_{4}\simeq(1,\{k_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
$\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{S}(\textit{tx}_{1})=\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{S}(\textit{tx}_{2})$ $=\frac{0+1+1+1+1+1}{6}$
$\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{S}(\textit{tx}_{3})$ $=\frac{2+1}{2}.$
Thus, $\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{S}(\textit{tx}_{1})+\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{S}(\textit{tx}_{2})=\frac{10}{6}>\frac{3}{2}=\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{S}(\textit{tx}_{3})$, which violates Transaction Bundling.
See 5.6
Proof C.2.
Consider a set of transactions $T$ and three transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\notin T$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$. Then, for any scheduler satisfying property (S2), we have:
$\textsf{gas}^{B}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}(\textit{tx}_{1})$ $+\textsf{gas}^{B}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}(\textit{tx}_{2})$
$=\frac{1}{2^{|T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}|-1}}\sum_{S\subseteq(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})\setminus\{\textit{tx}_{1}\}}\left[v(S\cup\{\textit{tx}_{1}\})-v(S)\right]$
$+\frac{1}{2^{|T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}|-1}}\sum_{S\subseteq(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})\setminus\{\textit{tx}_{2}\}}\left[v(S\cup\{\textit{tx}_{2}\})-v(S)\right]$
$=\frac{1}{2^{|T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}|-1}}\sum_{S\subseteq T}\left[v(S\cup\{\textit{tx}_{1}\})-v(S)+v(S\cup\{\textit{tx}_{1},\textit{tx}_{2}\})-v(S\cup\{\textit{tx}_{2}\})\right]$
$+\frac{1}{2^{|T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}|-1}}\sum_{S\subseteq T}\left[v(S\cup\{\textit{tx}_{2}\})-v(S)+v(S\cup\{\textit{tx}_{1},\textit{tx}_{2}\})-v(S\cup\{\textit{tx}_{1}\})\right]$
$=\frac{2}{2^{|T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}|-1}}\sum_{S\subseteq T}\left[v(S\cup\{\textit{tx}_{1},\textit{tx}_{2}\})-v(S)\right]$
$\leq\frac{1}{2^{|T\cup\{\textit{tx}_{3}\}|-1}}\sum_{S\subseteq T}\left[v(S\cup\{\textit{tx}_{3}\})-v(S)\right]\quad=\quad\textsf{gas}^{B}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).$
See 5.7
Proof C.3.
Consider a set of transactions $T$ and three transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\notin T$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$. Then, for any scheduler satisfying property (S2), we have:
$\textsf{gas}^{\textit{TPM}}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}(\textit{tx}_{1})$ $+\textsf{gas}^{\textit{TPM}}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}(\textit{tx}_{2})$
$=\frac{t(\textit{tx}_{1})}{\sum_{\textit{tx}^{\prime}\in T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}t(\textit{tx}^{\prime})}\cdot v(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})+\frac{t(\textit{tx}_{2})}{\sum_{\textit{tx}^{\prime}\in T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}t(\textit{tx}^{\prime})}\cdot v(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})$
$=\frac{t(\textit{tx}_{1})+t(\textit{tx}_{2})}{\sum_{\textit{tx}^{\prime}\in T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}t(\textit{tx}^{\prime})}\cdot v(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})$
$=\frac{t(\textit{tx}_{3})}{\sum_{\textit{tx}^{\prime}\in T\cup\{\textit{tx}_{3}\}}t(\textit{tx}^{\prime})}\cdot v(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})$
$\leq\frac{t(\textit{tx}_{3})}{\sum_{\textit{tx}^{\prime}\in T\cup\{\textit{tx}_{3}\}}t(\textit{tx}^{\prime})}\cdot v(T\cup\{\textit{tx}_{3}\})\quad=\quad\textsf{gas}^{\textit{TPM}}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).$
See 5.8
Proof C.4.
Consider the set of transactions $T=\{\textit{tx}_{4}\}$ and three transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\notin T$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$, where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(1,\{k_{1}\})$, $\textit{tx}_{3}\simeq(2,\{k_{1}\})$, and $\textit{tx}_{4}\simeq(1,\{k_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
$\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{\textit{ESM}}(\textit{tx}_{1})=\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{\textit{ESM}}(\textit{tx}_{2})=\frac{3}{3},\quad\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{\textit{ESM}}(\textit{tx}_{3})=\frac{3}{2}.$
Thus, $\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{\textit{ESM}}(\textit{tx}_{1})+\textsf{gas}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}^{\textit{ESM}}(\textit{tx}_{2})=2>\frac{3}{2}=\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}^{\textit{ESM}}(\textit{tx}_{3})$, which violates Transaction Bundling.
See 5.9
Proof C.5.
Consider a set of transactions $T$ and three transactions $\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\notin T$ such that $\textit{tx}_{3}$ is the concatenation of $\textit{tx}_{1}$ and $\textit{tx}_{2}$. Then, for any scheduler satisfying property (S2), we have:
$\textsf{gas}^{\textit{XSM}}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}(\textit{tx}_{1})$ $+\textsf{gas}^{\textit{XSM}}_{T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}}(\textit{tx}_{2})$
$=\frac{v(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})}{3^{|T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}|}}+\frac{v(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})}{3^{|T\cup\{\textit{tx}_{1},\textit{tx}_{2}\}|}}$
$=\frac{2}{3}\cdot\frac{v(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})}{3^{|T\cup\{\textit{tx}_{3}\}|}}$
$<\frac{v(T\cup\{\textit{tx}_{1},\textit{tx}_{2}\})}{3^{|T\cup\{\textit{tx}_{3}\}|}}$
$\leq\frac{v(T\cup\{\textit{tx}_{3}\})}{3^{|T\cup\{\textit{tx}_{3}\}|}}=\textsf{gas}^{\textit{XSM}}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3}).$
C.2 Set Inclusion (Property˜4)
See 5.10
Proof C.6.
Consider the transaction sets $T=\{\textit{tx}_{1}\}$, $T_{1}=\{\textit{tx}_{2},\textit{tx}_{3}\}$, and $T_{2}=\{\textit{tx}_{2},\textit{tx}_{3},\textit{tx}_{4}\}$ where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(1,\{k_{2}\})$, $\textit{tx}_{3}\simeq(1,\{k_{2}\})$, and $\textit{tx}_{4}\simeq(1,\{k_{1}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
$\textsf{gas}_{T\cup T_{1}}^{S}(T_{1})=\frac{5}{6}+\frac{5}{6},\quad\quad\textsf{gas}_{T\cup T_{2}}^{S}(T_{2})=\frac{12}{24}+\frac{12}{24}+\frac{12}{24}.$
Thus, $\textsf{gas}_{T\cup T_{1}}^{S}(T_{1})=\frac{10}{6}>\frac{36}{24}=\textsf{gas}_{T\cup T_{2}}^{S}(T_{2})$, which violates Set Inclusion.
See 5.11
Proof C.7.
Consider the transaction sets $T=\emptyset$, $T_{1}=\{\textit{tx}_{1},\textit{tx}_{2}\}$, and $T_{2}=\{\textit{tx}_{1},\textit{tx}_{2},\textit{tx}_{3}\}$ where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$, $\textit{tx}_{2}\simeq(1,\{k_{1}\})$, and $\textit{tx}_{3}\simeq(1,\{k_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
$\textsf{gas}_{T\cup T_{1}}^{B}(T_{1})=\frac{1+1}{2}+\frac{1+1}{2},\quad\quad\textsf{gas}_{T\cup T_{2}}^{B}(T_{2})=\frac{1+1+0+1}{4}+\frac{1+1+0+1}{4}+\frac{1+0+0+0}{4}.$
Thus, $\textsf{gas}_{T\cup T_{1}}^{B}(T_{1})=2>\frac{7}{4}=\textsf{gas}_{T\cup T_{2}}^{B}(T_{2})$, which violates Set Inclusion.
See 5.12
Proof C.8.
Consider a set of transactions $T$ and two sets of transactions $T_{1}\subseteq T_{2}$, disjoint from $T$. Then, for any scheduler satisfying property (S1), we have:
$\textsf{gas}^{\textit{TPM}}_{T\cup T_{1}}(T_{1})$ $=\frac{\sum_{\textit{tx}\in T_{1}}t(\textit{tx})}{\sum_{\textit{tx}^{\prime}\in T\cup T_{1}}t(\textit{tx}^{\prime})}\cdot v(T\cup T_{1})$
$\leq\frac{\sum_{\textit{tx}\in T_{1}}t(\textit{tx})+\sum_{\textit{tx}\in T_{2}\backslash T_{1}}t(\textit{tx})}{\sum_{\textit{tx}^{\prime}\in T\cup T_{1}}t(\textit{tx}^{\prime})+\sum_{\textit{tx}^{\prime}\in T_{2}\backslash T_{1}}t(\textit{tx}^{\prime})}\cdot v(T\cup T_{1})$
$=\frac{\sum_{\textit{tx}\in T_{2}}t(\textit{tx})}{\sum_{\textit{tx}^{\prime}\in T\cup T_{2}}t(\textit{tx}^{\prime})}\cdot v(T\cup T_{1})$
$\leq\frac{\sum_{\textit{tx}\in T_{2}}t(\textit{tx})}{\sum_{\textit{tx}^{\prime}\in T\cup T_{2}}t(\textit{tx}^{\prime})}\cdot v(T\cup T_{2})\quad=\quad\textsf{gas}^{\textit{TPM}}_{T\cup T_{2}}(T_{2}).$
See 5.13
Proof C.9.
Consider a set of transactions $T$ and two sets of transactions $T_{1}\subseteq T_{2}$, disjoint from $T$. Then, for any scheduler satisfying property (S1), we have:
$\textsf{gas}^{\textit{ESM}}_{T\cup T_{1}}(T_{1})$ $=\sum_{\textit{tx}\in T_{1}}\frac{v(T\cup T_{1})}{|T\cup T_{1}|}$
$=\frac{|T_{1}|}{|T\cup T_{1}|}\cdot v(T\cup T_{1})$
$\leq\frac{|T_{1}|+|T_{2}\backslash T_{1}|}{|T\cup T_{1}|+|T_{2}\backslash T_{1}|}\cdot v(T\cup T_{1})$
$=\frac{|T_{2}|}{|T\cup T_{2}|}\cdot v(T\cup T_{1})$
$\leq\frac{|T_{2}|}{|T\cup T_{2}|}\cdot v(T\cup T_{2})\quad=\quad\textsf{gas}^{\textit{ESM}}_{T\cup T_{2}}(T_{2}).$
See 5.14
Proof C.10.
Consider the transaction sets $T=\emptyset$, $T_{1}=\{\textit{tx}_{1}\}$, and $T_{2}=\{\textit{tx}_{1},\textit{tx}_{2}\}$, where $\textit{tx}_{1}\simeq(1,\{k_{1}\})$ and $\textit{tx}_{2}\simeq(1,\{k_{2}\})$. Then, for any number of threads $n\geq 2$ and the optimal scheduler, we have:
$\textsf{gas}_{T\cup T_{1}}^{\textit{XSM}}(T_{1})=\frac{1}{3^{1}},\quad\quad\textsf{gas}_{T\cup T_{2}}^{\textit{XSM}}(T_{2})=\frac{1}{3^{2}}+\frac{1}{3^{2}}.$
Thus, $\textsf{gas}_{T\cup T_{1}}^{\textit{XSM}}(T_{1})=\frac{1}{3}>\frac{2}{9}=\textsf{gas}_{T\cup T_{2}}^{\textit{XSM}}(T_{2})$, which violates Set Inclusion.
C.3 Storage Key-Time Monotonicity (Property˜3) and Strict Time Monotonicity (Property˜2)
See 5.15
Proof C.11.
Consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$. Then, for any scheduler satisfying property (S3):
$\textsf{gas}^{\textit{TPM}}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})$ $=\frac{t_{1}}{\sum_{\textit{tx}^{\prime}\in T\cup\{\textit{tx}_{1}\}}t(\textit{tx}^{\prime})}\cdot v(T\cup\{\textit{tx}_{1}\})$
$=\frac{t_{1}}{\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})+t_{1}}\cdot v(T\cup\{\textit{tx}_{1}\})$
$\leq\frac{t_{2}}{\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})+t_{2}}\cdot v(T\cup\{\textit{tx}_{1}\})$
$\leq\frac{t_{2}}{\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})+t_{2}}\cdot v(T\cup\{\textit{tx}_{2}\})\quad=\quad\textsf{gas}^{\textit{TPM}}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$
This establishes Storage Key-Time Monotonicity.
To also get strict Time Monotonicity, assume $t_{1}<t_{2}$ and $K_{1}=K_{2}$ in the above proof. Note that, when $\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})>0$, the function $\frac{t}{\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})+t}$ is strictly increasing in $t$, guaranteeing that the first inequality above is strict, so $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}^{\textit{TPM}}(\textit{tx}_{1})<\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}^{\textit{TPM}}(\textit{tx}_{2})$, as required. In the degenerate case where $\sum_{\textit{tx}^{\prime}\in T}t(\textit{tx}^{\prime})=0$, the fraction becomes 1 in both cases, but this can only happen when $T=\varnothing$, in which case the value $v$ increases strictly because $v(\{\textit{tx}_{1}\})=t_{1}<t_{2}=v(\{\textit{tx}_{2}\})$. Overall, the first inequality above again holds strictly, ensuring that $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}^{\textit{TPM}}(\textit{tx}_{1})<\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}^{\textit{TPM}}(\textit{tx}_{2})$ still holds, as required. Thus, the strict Time Monotonicity property is satisfied.